# Exploring Unseen Environments with Robots using Large Language and Vision Models through a Procedurally Generated 3D Scene Representation

### v1

@@ III-E Short term memory(STM) and reasoning
The agent then tasks an LVLM (LLaVA [35, 36]) to verify whether the detected segment in all the views is the target object or not. The determination of whether the detected segment represents the target object $G_{o}$ relies on the percentage of frames (different views) in which the LVLM (LLaVA) confirms the presence of the target object, serving as the decisive criterion for the final decision. If none of the detected segments are regarded as the target object then the current object selected for exploring is set to explored, the short term memory is cleared and the agent’s $action$ is set to $<explore\textunderscore scene>$.
III-F Execution level planner
@@
We make use of a goal oriented and episodic semantic mapping module [19], to build an episodic 2D scene obstacle map, frontier map and a goal map. In the phase where $action$ is $<explore\textunderscore scene>$, the goal map is same as that of the frontier map and when the $action$ is $<explore\textunderscore obj>$, the goal map is constructed in such a way that the robot can navigate to a region close to the object selected by the LLM. The goal map and the obstacle map is then parsed to a fast marching method planner, which plans execution level actionable sub-goals to reach the goal.
IV EXPERIMENT AND RESULTS
@@
We evaluate our stack to explore for a target object by building a 3D scene representation of the environment on the HomeRobot: Open Vocabulary Mobile Manipulation [37] simulation benchmark. This simulation benchmark have multi-room realistic environments, with a diverse and complex set of receptacles (eg: chair, table, couch, bed, toilet etc.) and articles (eg: pen, book, cellphone, tray, apple etc.) that can be manipulated.The environments, designed in Habitat [38, 39, 40] provide a cluttered and navigable scene having multiple rooms with manipulable objects placed on top of receptacles, to test various benchmarks related to navigation and manipulation of objects in unfamiliar environments. We chose this dataset over conventional object-nav datasets and benchmarks, because of the availability of a wide range of small objects, that are manipulable. The main focus of our work is to explore for these manipulable objects. However, the dataset provided by this benchmark does not consider the general semantics of the world while spawning objects on top of receptacles. For example, there is an episode in this dataset that spawns in an apple on top of a toilet. We went through the dataset and curated custom episodes that follow these semantics to test the stack.
@@
We test our stack for exploration by spawning the robot in a random location in the scene and asking it to find the target object within 500 steps. We evaluate the agents performance in navigating to the target object by considering success rate (SR) and success rate weighted by path length (SPL) which indicates the efficiency of the planner. The focus of our approach is in creating a planning pipeline that mimics human cognition in the way we perceive, store, express and regulate information. To test the efficacy of the approach compared to human level cognition, we tasked a set of random human volunteers to find the target object, for the same set of episodes that were given to the LLM based planner and measured the success rate (SR) and the success rate weighted by path length (SPL). The results of the experiments involving human and LLM agent is given in table I.
TABLE I: Experimental results
@@
TABLE II: Ablation analysis
@@
GPT-4 with STM and pruner 0.4375 0.2721472463
GPT-4 without STM 0.125 0.089943
GPT-4 withot pruner 0.182 0.1131
GPT-4 withot object captions 0.36363 0.2254
V DISCUSSION
V-A Comparison with human cognition
Experimental results from table I, shows the comparison between human agent and LLM agent in completing an object goal navigation task. Although GPT-4 based agent’s performance is far from human level performance, it has exhibited a similar thought process in exploring near certain objects in the scene, for finding the target object $G_{o}$. GPT-3.5 on the other hand failed to reason in most of the episodes, leading to an inefficient and meaningless exploration of the scene. Most of the failure scenarios with GPT-3.5 was when it failed to find the target object within the 500 step limit.
V-B Ablation studies
To signify the importance of having a short term memory based inference module, object captioning module and a detection pruner, we performed ablation analysis with these, the results of which are presented in table II.
@@ V-B Ablation studies
LLM based pruner is tasked to prune out smaller objects in an environment like pen, book, pillow etc, and keep object like chair, kitchen table, bed etc, which are important in understanding the semantics of the scene. By not doing so, the agent is flooded with a lot of redundant information, which can lead to inefficient plans. Table II shows the performance of the agent without the pruner. The agent without a pruning module engages in a lot of inefficient exploration, which can be seen with the reduced SPL, when compared to an agent with pruner.
@@ V-C Limitations
VI CONCLUSION
In this paper we introduced a framework to tackle the object goal navigation problem, by leveraging the generalization and semantic reasoning capabilities of Large Language Models and Large Vision Language Models to generate intelligent plans, by utilizing an efficient 3D scene representation of the environment. The agent embodied with this framework, was able to reason about its surrounding environment in a similar way to that of humans. Ablations studies shows the importance of having a short term memory (STM) based inference module, a pruner and semantic captions, which is in line with our architectural design inspired from human cognition. Future works may also explore the use of a more efficient representation, which the agent can use to reason about the environment.

### v2

@@ 3.5 Short term memory(STM) and reasoning
The agent then tasks an LVLM (LLaVA Liu et al. [36, 37]) to verify whether the detected segment in all the views is the target object or not. The determination of whether the detected segment represents the target object $G_{o}$ relies on the percentage of frames (different views) in which the LVLM (LLaVA) confirms the presence of the target object, serving as the decisive criterion for the final decision. If none of the detected segments are regarded as the target object then the current object selected for exploring is set to explored, the short term memory is cleared and the agent’s $action$ is set to $<explore\textunderscore scene>$.
3.6 Execution level planner
@@
We make use of a goal oriented and episodic semantic mapping module Chaplot et al. [22], to build an episodic 2D scene obstacle map, frontier map and a goal map. In the phase where $action$ is $<explore\textunderscore scene>$, the goal map is same as that of the frontier map and when the $action$ is $<explore\textunderscore obj>$, the goal map is constructed in such a way that the robot can navigate to a region close to the object selected by the LLM. The goal map and the obstacle map is then parsed to a fast marching method planner, which plans execution level actionable sub-goals to reach the goal.
4 EXPERIMENT AND RESULTS
@@
We evaluate our stack to explore for a target object by building a 3D scene representation of the environment on the HomeRobot: Open Vocabulary Mobile Manipulation Yenamandra et al. [38] simulation benchmark. This simulation benchmark have multi-room realistic environments, with a diverse and complex set of receptacles (eg: chair, table, couch, bed, toilet etc.) and articles (eg: pen, book, cellphone, tray, apple etc.) that can be manipulated.The environments, designed in Habitat Puig et al. [39], Szot et al. [40], Savva et al. [41] provide a cluttered and navigable scene having multiple rooms with manipulable objects placed on top of receptacles, to test various benchmarks related to navigation and manipulation of objects in unfamiliar environments. We chose this dataset over conventional object-nav datasets and benchmarks, because of the availability of a wide range of small objects, that are manipulable. The main focus of our work is to explore for these manipulable objects. However, the dataset provided by this benchmark does not consider the general semantics of the world while spawning objects on top of receptacles. For example, there is an episode in this dataset that spawns in an apple on top of a toilet. We went through the dataset and curated custom episodes that follow these semantics to test the stack.
@@
We test our stack for exploration by spawning the robot in a random location in the scene and asking it to find the target object within 500 steps. We evaluate the agents performance in navigating to the target object by considering success rate (SR) and success rate weighted by path length (SPL) which indicates the efficiency of the planner. The focus of our approach is in creating a planning pipeline that mimics human cognition in the way we perceive, store, express and regulate information. To test the efficacy of the approach compared to human level cognition, we tasked a set of random human volunteers to find the target object, for the same set of episodes that were given to the LLM based planner and measured the success rate (SR) and the success rate weighted by path length (SPL). The results of the experiments involving human and LLM agent is given in table 2.
Table 1: Experimental results
@@
Table 2: Ablation analysis
@@
GPT-4 0.437 0.272
GPT-4 w/o STM 0.125 0.089
GPT-4 w/o pruner 0.182 0.113
GPT-4 w/o captions 0.364 0.225
5 DISCUSSION
5.1 Comparison with human cognition
Experimental results from table 2, shows the comparison between human agent and LLM agent in completing an object goal navigation task. Although GPT-4 based agent’s performance is far from human level performance, it has exhibited a similar thought process in exploring near certain objects in the scene, for finding the target object $G_{o}$. GPT-3.5 on the other hand failed to reason in most of the episodes, leading to an inefficient and meaningless exploration of the scene. Most of the failure scenarios with GPT-3.5 was when it failed to find the target object within the 500 step limit.
5.2 Ablation studies
To signify the importance of having a short term memory based inference module, object captioning module and a detection pruner, we performed ablation analysis with these, the results of which are presented in table 2.
@@ 5.2 Ablation studies
LLM based pruner is tasked to prune out smaller objects in an environment like pen, book, pillow etc, and keep object like chair, kitchen table, bed etc, which are important in understanding the semantics of the scene. By not doing so, the agent is flooded with a lot of redundant information, which can lead to inefficient plans. Table 2 shows the performance of the agent without the pruner. The agent without a pruning module engages in a lot of inefficient exploration, which can be seen with the reduced SPL, when compared to an agent with pruner.
@@ 5.3 Limitations
6 CONCLUSION
In this paper we introduced a framework to tackle the object goal navigation problem, by leveraging the generalization and semantic reasoning capabilities of Large Language Models and Large Vision Language Models to generate intelligent plans, by utilizing an efficient 3D scene representation of the environment. Future works may also explore the use of a more efficient representation, which the agent can use to reason about the environment.
Acknowledgments
This work was supported by DAAD PPP-India program under project number 57622912. This work was supported by Ministry of Culture and Science of the State of North Rhine-Westphalia, Germany, through the KI-Starter funding program.